# ReX: Causal Discovery based on Machine Learning and Explainability techniques

### v1

@@ 5 Results with Single-cell protein network data
The results in Table 2 demonstrate that ReX significantly outperforms other methods across most evaluation metrics on this dataset, achieving the highest score in precision and F1 score, indicating both high specificity and a strong balance between precision and recall. While its recall of $0.471$ is moderate, it is still competitive and notably higher than that of most methods, except for CAM, which achieved a higher recall of $0.6$. In terms of structural accuracy, ReX also reports the lowest SHD (median of $9$), implying fewer structural errors compared to others, with FCI and CAM following at $13$ and $39$, respectively. In addition, ReX has a relatively good value for SID (median of $39$), indicating also a noteworthy intervention-based structure interpretation capability. Overall, ReX demonstrates strong precision and structural fidelity, supporting its robustness and reliability for causal discovery in this context.
@@ 6 Practical Considerations and Future Work
6.1 Explainability techniques
During method development, we compared Shapley values against other explainability techniques, such as feature importance, partial dependency plots, and LIME, and found that Shapley values outperformed all of them by a large margin. These techniques produced noisy results with limited practical application. As discussed in Section 2.1, while Shapley values may also provide insights into causal relationships, since they are not inherently causal, their application to obtain plausible causal graphs must be used in conjunction with causal inference techniques to explore and validate potential causal dynamics within the data.
@@ 6.2 Multicollinearity considerations
The implementation of Shapley values in SHAP by [4] includes model-agnostic explainers like GradientExplainer or TreeExplainer as a potential solution [36] to mitigate the effects of multicollinearity. Unlike other explainers that might treat features collectively or consider their joint distributions, interventional explainers operate under an interventional framework where the impact of each feature on the model’s output is computed independently, assuming other features remain constant. This approach aligns with the concept of partial derivatives, focusing on the sensitivity of the output to changes in each feature, treated in isolation. By doing so, interventional explainers effectively disentangle the intertwined influences of correlated features, making it easier to interpret the model in terms of causal relationships. This is particularly useful in deep learning models where the intricate interactions of features can be complex and opaque.
@@ 6.3 Future Work
Another research area for ReX is its application to datasets with distributional shifts, where the data-generating process may vary across different populations or over time. Recent work [39] has addressed this challenge by identifying features that remain causally invariant across changing distributions. Integrating such principles into the ReX could enhance its applicability to real-world datasets where the assumption of consistent distributions does not hold, improving its robustness and generalization capability.
@@ 7 Conclusion
Our experiments on five families of synthetic datasets and the Sachs single-cell protein-signaling dataset [32] illustrate the effectiveness and robustness of ReX. Not only does it recover causal relationships with high precision and low orientation error, but it also compares favorably with existing state-of-the-art methods. By providing interpretable insights into the underlying causal dynamics, ReX enables contrasting inferred causal links with domain expertise.
@@ Acknowledgements
Appendix A Parents selection example
@@
Fig. 7 illustrates how the parents selection described in Algorithm 2 mechanism works. On panel (a) there is a significant shift between the features that are more influential ($V1$ and $V2$) and the rest. On panel (${b}$), such shift is not so clear, though the first group of three ($V0,V2$ and $V4$) present a more pronounced influence than the rest. On top of each plot, the variables selected by Algorithm 2 are written to the right of the target variable, as parents ($\leftarrow$) of the target.
@@
In some cases, the proposed Algorithm 2 might discard features that are influential on the outcome of the regression of the target variable. This would be the case for example if, in Fig.7b, only the first variable $V0$ is selected instead of the first three variables (i.e., $V0$, $V2$ and $V4$). In those cases, manual supervision or special parametrization of the method would be required to ensure that the selected features are also the most influential. Indeed, an additional parametrization of Algorithm 2 allows to run the method in greedy mode, to return not only the features with higher impact but all whose value is above a given percentile of the distribution of SHAP values. This tuning can be automatically activated in those cases where the feature selection process leads to a plausible causal graph (Section 3.2.3) with a number of edges significantly below the max number of edges expected in a fully connected DAG, given by $\frac{p(p-1)}{2}$. We note that none of the experimental results described in Section 4 required the use of this or any other special parametrization of the main ReX method.
Appendix B SHAP discrepancy example

### v2

@@ 5.1 Single-cell protein network data
The results in Table 2 demonstrate that ReX significantly outperforms other methods across most evaluation metrics on this dataset, achieving the highest score in precision and F1 score, indicating both high specificity and a strong balance between precision and recall. While its recall of $0.471$ is moderate, it is still competitive and notably higher than that of most methods, except for CAM, which achieved a higher recall of $0.6$. In terms of structural accuracy, ReX also reports the lowest SHD (median of $9$), implying fewer structural errors compared to others, with FCI and CAM following at $13$ and $39$, respectively. In addition, ReX has a relatively good value for SID (median of $39$), indicating also a noteworthy intervention-based structure interpretation capability. Overall, ReX demonstrates strong precision and structural fidelity, supporting its robustness and reliability for causal discovery in this context (see Figure 7).
Figure 7: DAGs obtained by the rest of comparable methods on the Sachs datasets.
@@
5.2 Financial decision-making data
The ReX method was further evaluated on a synthetic financial dataset [34] designed to emulate real-world financial decision-making scenarios. This dataset captures the structural complexities typically encountered in financial credit risk assessments, including heterogeneous treatment effects, non-linear interactions, and positivity assumption violations. The dataset consists of 12 variables, with an observational sample of 12,000 entries. The variables include financial indicators such as credit history, debt levels, number of loans, and external debt exposure. The treatment variable represents the percentage of debt write-off assigned to customers, while the outcome variable reflects repayment probability. The known true causal DAG (see Fig. 8c) underlying the data generation process serves as ground truth for evaluating the effectiveness of different causal discovery methods (see [34] for details on the data generation process).
The causal discovery methods were applied ten times to different data subset samples (made of 1,000 samples), and performance was assessed based on precision, recall, F1-score, SHD and SID. ReX was applied following the same pipeline as in previous experiments. As with the single-cell protein network data, a hyperparameter search (300 iterations) was conducted to optimize the regressors used for SHAP value computation. Each dataset was split into training and validation subsets, with the causal discovery process applied to the entire dataset after model training. Bootstrapping (10 iterations) was employed to enhance stability in the inferred causal structure, for the ReX case. Two sample DAGs (best and worst) extracted from the bootstrapping process are show in Figure 8, in addition to the ground truth.
Table 3: Summary metrics for all considered methods on the Financial dataset.
Method Precision Recall F1 SHD SID
[…]
The summary metrics for all considered methods are provided in Table 3. The results indicate that ReX consistently outperforms other methods in terms of structural accuracy. With an F1-score of $0.850\pm 0.05$, ReX surpasses the second-best performing method, GES, which achieves an F1-score of $0.610\pm 0.06$. This improvement can be attributed to ReX ’s ability to effectively leverage machine learning explainability techniques for causal inference.
In terms of precision, ReX achieves a score of $0.915\pm 0.12$, which is only slightly lower than the highest-performing method in this regard, FCI, which reaches $0.967\pm 0.07$. However, ReX excels in recall with a significantly higher score of $0.799\pm 0.03$, substantially outperforming FCI and PC, both of which yield recall values of $0.375\pm 0.12$ and $0.375\pm 0.10$, respectively. This suggests that ReX is more effective at identifying true causal relationships without omitting key connections.
Furthermore, ReX achieves the lowest Structural Hamming Distance (SHD) and Structural Intervention Distance (SID) among all tested methods, with values of $4.6\pm 1.82$ and $6.4\pm 1.52$, respectively. These results highlight the method’s ability to recover the true causal structure with minimal errors and to maintain robustness in intervention-based causal queries.
Compared to constraint-based methods such as PC and FCI, ReX shows superior performance in terms of both recall and SHD. The performance of NOTEARS and LiNGAM remains significantly lower, with F1-scores of $0.553\pm 0.03$ and $0.170\pm 0.09$, respectively. The CAM method performs the worst, with an F1-score of $0.171\pm 0.11$, an SHD of $35.4\pm 9.1$, and the highest SID value of $83.4\pm 15.04$, indicating that it struggles to correctly infer causal relationships in this domain.
Figure 8: Two samples of the DAGs inferred by ReX on the financial dataset: (a) the DAG with the best inferred structure, and (b) the DAG with the worst inferred structure, as well as (c) the true DAG.
@@ 6 Practical considerations and future work
6.1 Selection of Shapley values for feature importance assessment
A fundamental design decision in ReX involves selecting a reliable measure of feature contributions to predict a target variable, which directly guides parent selection (Algorithm 2). Among various methods, Shapley values [5] were chosen due to their theoretical rigor and practical advantages in identifying stable and meaningful candidate parents.
Alternative methods have specific limitations. Model-based metrics, such as Gini importance, may bias results toward high-cardinality features or produce misleading attributions when features are correlated [35]. Permutation importance, although intuitive, struggles similarly in correlated scenarios, either underestimating collective feature importance or producing unrealistic permutations [36, 37]. Local explanation methods like LIME [38] require extensive aggregation for global insights and are sensitive to the choice of local neighborhoods and surrogate models [39].
Conversely, Shapley values as implemented through SHAP [4] provide several distinct benefits for ReX: (i) they uniquely satisfy key axiomatic properties (efficiency, symmetry, and additivity) that enable principled attribution [5, 6]; (ii) they inherently accommodate feature interactions and redundancies by averaging contributions across all possible feature coalitions, ensuring stable attribution even in scenarios of high multicollinearity, as demonstrated empirically (Appendix A); (iii) they exhibit a theoretical and empirically validated link to conditional dependence (Section 2.2, Appendix A), aligning closely with the requirements of causal discovery under faithfulness; and (iv) they benefit from robust, model-specific implementations (e.g., TreeExplainer for GBTs, GradientExplainer for DFNs), improving computational efficiency and reliability [4, 7].
Due to their theoretical soundness, robustness to feature correlations, and clear interpretability regarding conditional dependence, Shapley values are particularly well-suited to identifying potential causal parents in ReX.
@@ 6.2 Multicollinearity considerations
The implementation of Shapley values in SHAP by [4] includes model-agnostic explainers like GradientExplainer or TreeExplainer as a potential solution [7] to mitigate the effects of multicollinearity.
Specifically, these “interventional” variants of SHAP explainers, as discussed by Chen et al. [7] in contrast to “observational” approaches that use conditional expectations, aim to estimate a feature’s effect by breaking its correlations with other features not currently in the evaluated coalition. When computing the expected model output for a subset of features $S$, features $X_{k}\notin S$ are effectively marginalized out by sampling from their empirical distribution (often a background dataset) independently of the features in $S$, rather than conditioning on them.
For instance, TreeExplainer, when applied to tree-based models like GBTs, can implement this through path-dependent feature perturbation algorithms that effectively average over the marginal distribution of “absent” features encountered along decision paths. GradientExplainer, used for DFNs, approximates SHAP values and typically relies on a background dataset to represent features not in the current coalition; if these are sampled independently, its behavior aligns with an interventional perspective.
Unlike other explainers that might treat features collectively or consider their joint distributions, interventional explainers operate under an interventional framework where the impact of each feature on the model’s output is computed independently, assuming other features remain constant. This approach aligns with the concept of partial derivatives, focusing on the sensitivity of the output to changes in each feature, treated in isolation. By doing so, interventional explainers effectively disentangle the intertwined influences of correlated features, making it easier to interpret the model in terms of causal relationships. This is particularly useful in deep learning models where the intricate interactions of features can be complex and opaque.
@@
Indeed, our synthetic experiments (detailed in Appendix A) provide empirical evidence for this robustness. In scenarios with highly collinear predictor variables, the Shapley values computed remained stable across repeated repeated trials ($\sigma/\mu<5\%$ for $\phi$ across 50 replicates) and provided an intuitive partitioning of importance between the correlated features, a task that can be challenging for other interpretation methods or even for establishing stable DAG structures using iterative CI-tests.
6.3 Stability of the discovered DAG
The stability of the causal graph discovered by ReX is primarily addressed through the bootstrapping mechanism (Algorithm 1) and the subsequent frequency-based thresholding using the $\tau$ parameter (Section 3.2.3). An edge is incorporated into the initial graph for each regressor type ($G_{DFN}$ or $G_{GBT}$) only if its selection frequency across $T$ bootstrap iterations surpasses $\tau$. This procedure ensures that the foundational structure is based on relationships that consistently emerge despite resampling of the data, thereby filtering out less robust, potentially spurious connections.
While ReX does not compute a separate, global quantitative metric for DAG stability over multiple identical full runs, the design inherently promotes the stability of the identified edges. The consistency and relatively low variance observed in performance metrics across diverse synthetic datasets (e.g., F1 scores in Figure 14) further provide indirect evidence of the method’s ability to produce reliable DAGs. The overall stability of the final $G_{\text{R{e}X}}$ also depends on the robustness of the hyperparameter optimization, the ANM-based edge orientation (Section 3.3), and the cycle resolution mechanism (Section 3.4).
@@ 6.4 Future work
Another research area is its application to datasets with distributional shifts, where the data-generating process may vary across different populations or over time. Recent work [43] has addressed this challenge by identifying features that remain causally invariant across changing distributions. Integrating such principles into the ReX could enhance its applicability to real-world datasets where the assumption of consistent distributions does not hold, improving its robustness and generalization capability.
@@
While our SHAP discrepancy measure (Section 3.4) offers a novel, explainability-guided heuristic for resolving cycles that may arise from the union of DAGs, a comparative evaluation against other cycle-breaking strategies could be pursued. Future investigations might explore graph-theoretic approaches, such as heuristics for the Feedback Arc Set problem (e.g., [44]), or methods adapted from score-based Bayesian network learning to further refine this crucial step.
@@ 7 Conclusion
Our experiments on five families of synthetic datasets, the Sachs single-cell protein-signaling dataset [31], and the financial decision-making data [34] illustrate the effectiveness and robustness of ReX. Not only does it recover causal relationships with high precision and low orientation error, but it also compares favorably with existing state-of-the-art methods. By providing interpretable insights into the underlying causal dynamics, ReX enables contrasting inferred causal links with domain expertise.
@@ Acknowledgements
Appendix A Experimental validation of Shapley values versus statistical dependence
This appendix details the design, implementation, and results of a series of synthetic data experiments specifically conducted to elucidate the relationship between Shapley feature importance scores (as utilized in ReX) and traditional measures of statistical dependence (marginal and conditional). These experiments complement the broader performance benchmarks of ReX presented in Section 4 by providing a focused analysis on the behavior of Shapley values in canonical causal structures known to pose challenges for causal discovery algorithms that rely on CI tests. The goal is to empirically validate the theoretical foundations discussed in Section 2.2 and address common questions regarding the interpretation of Shapley values in causal contexts, particularly concerning confounders, mediators, colliders, and multicollinearity.
A.1 Experimental design and methodology
The experiment was designed to systematically compare Shapley values with statistical dependence measures across various controlled causal scenarios. The key steps are outlined below:
Four fundamental causal structures involving three variables ($X$, $Y$, $Z$, where $Y$ is the target variable for prediction) or two highly correlated parent variables ($X_{1},X_{2}$) were generated (see Table 4).
Structure Description
Confounder $Z\rightarrow X$, $Z\rightarrow Y$. This structure induces a spurious correlation between $X$ and $Y$ due to the common cause $Z$.
Chain (Mediator) $X\rightarrow Z\rightarrow Y$. Here, the influence of $X$ on $Y$ is mediated by $Z$.
Collider $X\rightarrow Z\leftarrow Y$. In this structure, $X$ and $Y$ are marginally independent but become conditionally dependent given their common effect $Z$.
Collinear Parents $X_{1}\rightarrow Y$, $X_{2}\rightarrow Y$, with $X_{1}\approx X_{2}$. This scenario tests behavior under high multicollinearity between parent nodes. For this structure, $Y$ is the target, and $X_{1},X_{2}$ are the features of interest.
Table 4: Causal structures considered in the analysis
For the Confounder, Chain, and Collider structures, the target variable for prediction and Shapley value computation is always $Y$. The features of interest for which Shapley values and dependence measures are reported are $X$ and $Z$.
Data was generated from linear Gaussian structural equations with coefficients set to 1.0, except in the collinear case, where a small perturbation induced multicollinearity. Each dataset comprised $n=5000$ samples with noise terms drawn from a $\mathcal{N}(0,0.10^{2})$ distribution. The process was repeated $R=50$ times per structure to ensure robust estimation of means and standard deviations for all metrics.
A gradient-boosted tree model (XGBoost) was fitted to predict the target variable $Y$ from the relevant features ($X,Z$ or $X_{1},X_{2}$, depending on the causal structure). The model hyperparameters were fixed at 200 trees, maximum depth of 4, and learning rate of 0.05 across all experiments to isolate the effects of data structure, despite the linearity of the data generation process.
Shapley values were computed using TreeExplainer from the SHAP library [4] to quantify feature importance in predicting $Y$. For each of the $R=50$ replicates, mean absolute Shapley values ($\phi$) per feature were calculated on a held-out test set of 1,000 points.
To compare Shapley values with traditional statistical measures, marginal and conditional dependence metrics were computed between each relevant feature and the target $Y$. Marginal dependence was quantified using the Pearson correlation coefficient $\rho(\text{Feature},Y)$. Conditional dependence was assessed via partial Pearson correlation coefficients, e.g., $\rho(X,Y\mid Z)$ for structures involving additional non-target features. Statistical significance of conditional dependence was evaluated through p-values calculated using Fisher’s Z-transform.
A.2 Aggregated results
The mean and standard deviation (s.d.) of the computed metrics across the $R=50$ replications are presented in Table 5. Columns are: $\overline{\rho}$ (mean marginal Pearson correlation with $Y$), $\sigma_{\rho}$ (s.d. of marginal Pearson correlation), $\overline{\rho_{p}}$ (mean partial correlation with $Y$, conditioned as described above), $\sigma_{\rho_{p}}$ (s.d. of partial correlation), $\overline{p}$ (mean $p$-value for the partial correlation), $\sigma_{p}$ (s.d. of $p$-value), $\overline{\phi}$ (mean absolute Shapley value for predicting $Y$), and $\sigma_{\phi}$ (s.d. of mean absolute Shapley value).
Table 5: Aggregated results (mean $\pm$ s.d. over 50 replications) comparing Shapley values with marginal and conditional dependence measures for different causal structures. The target variable is $Y$. For Confounder, Chain, and Collider, features are $X$ and $Z$. For Collinear, features are $X_{1}$ and $X_{2}$.
Structure Feature $\overline{\rho}$ $\sigma_{\rho}$ $\overline{\rho_{p}}$ $\sigma_{\rho_{p}}$ $\overline{p\text{-val}}$ $\sigma_{p\text{-val}}$ $\overline{|\phi|}$ $\sigma_{|\phi|}$
Confounder X 0.990 0.000 0.000 0.013 0.514 0.277 0.356 0.009
Z 0.995 0.000 0.706 0.007 0.000 0.000 0.447 0.011
Chain X 0.990 0.000 -0.001 0.011 0.625 0.293 0.426 0.009
Z 0.995 0.000 0.707 0.006 0.000 0.000 0.372 0.008
Collider X 0.000 0.014 -0.990 0.000 0.000 0.000 0.556 0.018
Z 0.705 0.008 0.995 0.000 0.000 0.000 0.873 0.023
Collinear $X_{1}$ 0.995 0.000 0.097 0.015 0.000 0.000 0.435 0.012
$X_{2}$ 0.995 0.000 0.102 0.015 0.000 0.000 0.361 0.010
A.3 Interpretation and discussion of results
The results from Table 5 provide quantitative evidence supporting the theoretical connection between Shapley values and conditional dependence, as articulated in Section 2.2 (particularly Equation (11)).
A.3.1 Divergence from single-Set CI tests
The experiments show that Shapley values are not mere proxies for marginal correlation ($\rho$) nor for conditional independence tests based on a single, specific conditioning set (like the partial correlation $\rho_{p}$ and its $p$-value).
  • Confounder ($Z\rightarrow X,\ Z\rightarrow Y$): For feature $X$, the marginal correlation $\rho(X,Y)$ is very high (mean 0.990) due to the confounder $Z$. However, when conditioned on $Z$, the partial correlation $\rho_{p}(X,Y|Z)$ is effectively zero (mean 0.000), with a high mean $p$-value (0.514), indicating $X\mathbin{\perp\!\!\!\perp}Y|Z$. Despite this conditional independence given $Z$, the Shapley value for $X$ (mean $|\phi|_{X}=0.356$) is substantial. This occurs because, according to Equation (10), Shapley values consider all coalitions $\mathcal{S}$. In coalitions where $Z\notin\mathcal{S}$, $X$ still carries predictive information about $Y$ (as $X$ is a proxy for $Z$’s influence). The non-zero Shapley value reflects this average utility across all contexts. Feature $Z$, the true direct cause of $Y$ (in this simplified model), has a high partial correlation (mean $\rho_{p}(Z,Y|X)=0.706$, $p$-value $\approx 0$) and the highest Shapley value (mean $|\phi|_{Z}=0.447$).
  • Chain ($X\rightarrow Z\rightarrow Y$): Similarly for feature $X$, $\rho(X,Y)$ is high (mean 0.990). When conditioned on the mediator $Z$, $\rho_{p}(X,Y|Z)$ is effectively zero (mean -0.001), with a high mean $p$-value (0.625), indicating $X\mathbin{\perp\!\!\!\perp}Y|Z$. Yet, $|\phi|_{X}$ is substantial (mean 0.426). This is because in coalitions $\mathcal{S}$ where $Z\notin\mathcal{S}$, the path $X\rightarrow Z\rightarrow Y$ is ”open” in terms of information flow from $X$ to $Y$ for the predictive model, making $X$ valuable. The direct parent $Z$ has a high partial correlation $\rho_{p}(Z,Y|X)$ (mean 0.707, $p$-value $\approx 0$) and a significant Shapley value (mean $|\phi|_{Z}=0.372$).
These cases demonstrate that if a standard CI test (like Fisher-Z on partial correlation) indicates conditional independence given a specific variable (e.g., $Z$), the Shapley value for $X$ can still be high. This is because the Shapley value integrates over all possible conditioning sets (coalitions), capturing the feature’s predictive contribution in contexts where the specific variable $Z$ is not part of the conditioning set. This aligns with the theoretical premise that Shapley values offer a more global assessment of feature importance.
A.3.2 Collider structure and conditional dependence
  • Collider ($X\rightarrow Z\leftarrow Y$): Feature $X$ is marginally independent of $Y$ (mean $\rho(X,Y)=0.000$). However, when conditioned on the collider $Z$, $X$ and $Y$ become strongly dependent (mean $\rho_{p}(X,Y|Z)=-0.990$, $p$-value $\approx 0$). The Shapley value for $X$ (mean $|\phi|_{X}=0.556$) is high, reflecting this strong conditional dependence that emerges when $Z$ is in the coalition. Feature $Z$ itself, being a child of both $X$ and $Y$ (and thus highly informative about $Y$ when $X$ is known, and vice-versa for the model), has the highest Shapley value (mean $|\phi|_{Z}=0.873$). This scenario illustrates that Shapley values correctly identify the importance of variables whose relevance is only revealed upon conditioning.
A.3.3 Robustness and stability in high correlation scenarios
The Collinear Parents case addresses concerns about behavior with highly correlated variables.
  • Collinear Parents ($X_{1}\approx X_{2}\rightarrow Y$): Both $X_{1}$ and $X_{2}$ are highly correlated with $Y$ marginally (mean $\rho\approx 0.995$). The partial correlations $\rho_{p}(X_{1},Y|X_{2})$ and $\rho_{p}(X_{2},Y|X_{1})$ are low (means 0.097 and 0.102 respectively), indicating that once one collinear parent is known, the other offers little unique additional linear information. Despite these low partial correlations, their $p$-values are effectively zero due to the large sample size, indicating statistical significance. Crucially, the Shapley values for $X_{1}$ (mean $|\phi|_{X1}=0.435$) and $X_{2}$ (mean $|\phi|_{X2}=0.361$) are substantial and intuitively distribute the importance between the two highly correlated features. The standard deviations for these Shapley values are very small ($0.012$ and $0.010$ respectively), highlighting the stability of Shapley attributions in this scenario. This contrasts with the potential instability or difficulty in interpreting individual coefficients in a linear regression model facing severe multicollinearity. This empirical finding supports the discussion in Section 6.2 regarding the robustness of interventional SHAP explainers.
A.3.4 Confirmation of theoretical connection (Equation (10))
The results presented along this appendix confirm the theoretical connection established in 2.2, which can equivalently represented as:
$\phi_{j}\;=\;\bar{\delta}_{j}\times\Pr_{\mathcal{S}\sim w}\!\bigl[X_{j}\not\!\perp\!\!\!\perp Y\mid\mathcal{S}\bigr]$ (13)
[…]
  • When a feature $X_{j}$’s conditional dependence on $Y$ is only evident in specific contexts (e.g., $X$ in Confounder/Chain is dependent on $Y$ mainly when $Z$ is not in the coalition), its Shapley value is moderated but still non-zero, reflecting this averaged importance.
  • The low standard deviations of the mean Shapley values across 50 runs (typically $5\%$ of the mean, often much lower) underscore the stability of this importance measure under the experimental conditions.
A.4 Conclusion of experimental validation
This synthetic study empirically illustrates that Shapley values utilized in ReX capture aspects of feature importance beyond those measured by standard statistical dependence tests. Specifically, Shapley values demonstrate increased robustness under multicollinearity, reveal important distinctions from conventional measures, and substantiate their theoretical relationship to conditional dependence in aggregate. Collectively, these results support the integration of Shapley values into causal discovery frameworks such as ReX.
Appendix B Parents selection example
@@
Fig. 9 illustrates how the parents selection described in Algorithm 2 mechanism works. On panel (a) there is a significant shift between the features that are more influential ($V1$ and $V2$) and the rest. On panel (b{b}), such shift is not so clear, though the first group of three ($V0,V2$ and $V4$) present a more pronounced influence than the rest. On top of each plot, the variables selected by Algorithm 2 are written to the right of the target variable, as parents ($\leftarrow$) of the target.
@@
While Algorithm 2 is designed for fully automated parent selection, we acknowledge the theoretical possibility of edge cases where its standard DBSCAN-based clustering (detailed in § 3.2.2) might discard features that a user, upon closer inspection of SHAP value distributions (as exemplified in Figure 9), could still deem influential, particularly if SHAP values are very closely clustered or if the automated separation appears too conservative for a specific application. For instance, in a scenario like that depicted in Figure 9b, if the automated process were to hypothetically select only variable $V0$ instead of all three visually distinct top variables ($V0,V2,V4$) due to such an edge case in the clustering, a user might seek further options.
For such specific situations, where an application user might desire finer-grained control or wish to incorporate domain expertise, the concept of “manual supervision” is noted here as a practical recourse. This could involve, for example, a domain expert directly reviewing the SHAP value distributions and, based on their judgment, potentially adjusting the automatically selected parent set. As an alternative built-in option for users seeking a more inclusive selection without direct manual intervention, an additional parametrization of Algorithm 2 allows it to be run in a greedy mode. This mode can be configured to select all features whose SHAP values exceed a user-defined percentile of the distribution, rather than relying solely on the primary DBSCAN clustering outcome. Furthermore, as an optional heuristic, this greedy tuning can be automatically activated if the standard feature selection process results in a causal graph (see § 3.2.3) with a number of edges significantly below what might be expected for a densely connected system (e.g., compared to $\frac{p(p-1)}{2}$).
It is crucial to emphasize, however, that the performance and results reported for ReX throughout this paper (Section 4) are based on its fully automated execution, without any such manual supervision or the use of the ‘greedy mode’ or other special parametrizations of Algorithm 2. The discussion of these alternative operational modes is included in this appendix purely as a practical consideration for users who might apply ReX to new, potentially unique, or particularly challenging datasets where such adjustments could be exceptionally beneficial for fine-tuning the discovery process.
Appendix C SHAP discrepancy example